OpenStudy (anonymous):

30:42::110: ? a) 168 b) 144 c) 156 d) 132 e) 152

OpenStudy (jiteshmeghwal9):

\[30:42::110:x\]\[\frac{30}{42}=\frac{110}{x}\]By cross multiplication\[30 \times x=42 \times 110\]Solve this for \(x\).

OpenStudy (jiteshmeghwal9):

Gt it @kingkos ???

OpenStudy (anonymous):

154?

OpenStudy (jiteshmeghwal9):

right :)

OpenStudy (anonymous):

thanks

OpenStudy (jiteshmeghwal9):

yw :)